Determine the pH during the titration of 14.7 mL of 0.118 M hydroiodic acid by 8.94×10-2 M sodium hydroxide at the following points:

(1) Before the addition of any sodium hydroxide

(2) After the addition of 9.70 mL of sodium hydroxide

(3) At the equivalence point

(4) After adding 24.6 mL of sodium hydroxide
Answer #1

1)when 0.0 mL of NaOH is added

Given:

M(HI) = 0.118 M

V(HI) = 14.7 mL

M(NaOH) = 0.0894 M

V(NaOH) = 0 mL

mol(HI) = M(HI) * V(HI)

mol(HI) = 0.118 M * 14.7 mL = 1.7346 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.0894 M * 0 mL = 0 mmol

We have:

mol(HI) = 1.735 mmol

mol(NaOH) = 0 mmol

0 mmol of both will react

remaining mol of HI = 1.735 mmol

Total volume = 14.7 mL

[H+]= mol of acid remaining / volume

[H+] = 1.735 mmol/14.7 mL

= 0.118 M

use:

pH = -log [H+]

= -log (0.118)

= 0.9281

Answer: 0.928

2)when 9.7 mL of NaOH is added

Given:

M(HI) = 0.118 M

V(HI) = 14.7 mL

M(NaOH) = 0.0894 M

V(NaOH) = 9.7 mL

mol(HI) = M(HI) * V(HI)

mol(HI) = 0.118 M * 14.7 mL = 1.7346 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.0894 M * 9.7 mL = 0.8672 mmol

We have:

mol(HI) = 1.735 mmol

mol(NaOH) = 0.8672 mmol

0.8672 mmol of both will react

remaining mol of HI = 0.8674 mmol

Total volume = 24.4 mL

[H+]= mol of acid remaining / volume

[H+] = 0.8674 mmol/24.4 mL

= 3.555*10^-2 M

use:

pH = -log [H+]

= -log (3.555*10^-2)

= 1.4492

Answer: 1.45

3)

So, pH at equivalence point will be neutral which is 7.00

Answer: 7.00

4)when 24.6 mL of NaOH is added

Given:

M(HI) = 0.118 M

V(HI) = 14.7 mL

M(NaOH) = 0.0894 M

V(NaOH) = 24.6 mL

mol(HI) = M(HI) * V(HI)

mol(HI) = 0.118 M * 14.7 mL = 1.7346 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.0894 M * 24.6 mL = 2.1992 mmol

We have:

mol(HI) = 1.735 mmol

mol(NaOH) = 2.199 mmol

1.735 mmol of both will react

remaining mol of NaOH = 0.4646 mmol

Total volume = 39.3 mL

[OH-]= mol of base remaining / volume

[OH-] = 0.4646 mmol/39.3 mL

= 1.182*10^-2 M

use:

pOH = -log [OH-]

= -log (1.182*10^-2)

= 1.9273

use:

PH = 14 - pOH

= 14 - 1.9273

= 12.0727

Answer: 12.07